OpenStudy (anonymous):

Belle's Bake Shop makes croissants that cost $1.75 each. Past experience shows that 10 percent of croissants will spoil and have to be discarded. Assuming Belle wants a 45 percent markup based on cost and produces 300 croissants, each croissant should sell for

OpenStudy (compassionate):

I'm going to answer this. No one snipe me, please.

OpenStudy (anonymous):

:)

OpenStudy (compassionate):

$1.75*300=$525 cost 1.45=$761.25, 300*0.9=270 croissants sold, $761.25/270=$2.82 selling price per croissant.

OpenStudy (compassionate):

===================================> Best, please.